OpenStudy (anonymous):

Can someone help me. −4|3x + 7| = −14

OpenStudy (anonymous):

Don't just give the answer

OpenStudy (anonymous):

divide by \(-4\) first what do you get?

OpenStudy (anonymous):

3.5

OpenStudy (anonymous):

Help!!!!!

OpenStudy (anonymous):

you get \[|3x+7|=3.5\] right?

OpenStudy (anonymous):

then solve \[3x+7=3.5\] and also solve \[3x+7=-3.5\]

OpenStudy (anonymous):

so x would equal -1.166

OpenStudy (anonymous):

never mind I get it

OpenStudy (anonymous):

Thanks guys